what is 6 4/5 divided by 17?
what is 3 4/9 divided by 2 1/3?
what is 4 3/4 divided by 2 1/2?

Answer by checkley79(3341)

6 4/5=34/5
(34/5)/17=(34/5)*1/17=2/5 ans.
3 4/9=31/9
2 1/3=7/3
(31/9)/(7/3) invert & multiply
(31/9)/(3/7)=(31*3)/(9*7)=93/63 reduce by dividing by 21.
93/63=31/21 ans.